OpenStudy (anonymous):

Solve log3^3 + log3^x = 5. Is it: 3log3 + xlog3 = 5 3xlog3 = 5 x ~ 3,5 ? :/

OpenStudy (openstudier):

If you have two log with the same base and your adding them, that is the same thing as multiplying the terms together.

OpenStudy (openstudier):

\[\log _{a}b+\log_{a}c=\log_{a}bc\]

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school, I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador

OpenStudy (openstudier):

\[\log3^{3}+\log3^{x}=5\]

OpenStudy (openstudier):

\[\log3^{3x}=5\]

OpenStudy (openstudier):

\[3xlog3=5\]

OpenStudy (anonymous):

and how do I find x?

OpenStudy (openstudier):

\[x=\frac{ 5 }{ 3\log3 }\]

OpenStudy (raden):

3^x * 3^3 not equal 3^(3x) it is 3^(x+3), @OpenStudier

OpenStudy (raden):

well, i will use ur step above, @Haruhi

OpenStudy (raden):

log3^3 + log3^x = 5 3log3 + xlog3 = 5 xlog3 = 5 - 3log3 x = (5 - 3log3)/log3 it is, easy right ?

OpenStudy (anonymous):

Thanks so much! That helped a lot! So loga^x + loga^y is never = loga^xy? :/

OpenStudy (raden):

first, i agree about the property like openstudier said above : log_a b+log_a c =log_a (bc) but, in the property of exponent. if given a*x times a*y then it can be a^(x+y) not a^(xy)

OpenStudy (anonymous):

Thanks!! ^-^

OpenStudy (raden):

you're welcome :D